REVTEX 3.0 Released November 10, 1992

REVTEX Information For AIP Authors

Instructions to authors for preparing compuscripts to be submitted to AIP journals in REVTEX 3.0 format

Contents

I  Introduction  3-1
II  Participating Journals  3-1

III Where to turn for help 3-1

A Prior to manuscript acceptance …. 3-1 1 Procedural questions . . . . . . . . 3-1 2 REVTEX Questions . . . . . . . . . 3-2 B After Manuscript Acceptance . . . . . . 3-2 C AIP Contacts . . . . . . . . . . . . . . 3-2 D Editorial Office Contacts . . . . . . . . 3-2

IV Other packages that can be used with REVTEX 3-3

V  Basic Information  3-3

VI How To Use REVTEX 3-4

I. INTRODUCTION

The American Institute of Physics will be translating REVTEX files into code that can be used by our Xyvision composition system. Composition (page make-up) will take place using the Xyvision-coded file. For this reason, compuscripts will look exactly like conventionally processed articles when they appear in print, and authors should check proofs carefully.

After the REVTEX file is converted to Xyvision code, the article will follow conventional processing schedules. After sufficient experience with compuscript processing, the AIP will examine whether page-charge reductions and/or accelerated production schedules for compuscripts are possible.

II. PARTICIPATING JOURNALS

The American Institute of Physics publishes more than twenty journals at the time this Guide was drafted. The following journals are published by the American Institute of Physics and most are covered in this Guide.

• Astronomical Journal^{∗}

• American Journal of Physics

• Applied Physics Letters^{∗} • Chaos

• Computers in Physics

• Journal of Applied Physics

• The Journal of the Acoustical Society of America

• The Journal of Chemical Physics

• Journal of Mathematical Physics

• Journal of Rheology

• Journal of Vacuum Science and Technology A

• Journal of Vacuum Science and Technology B

• Medical Physics

• Publications of the Astronomical Society of the Pacific

• Physics of Fluids A

• Physics of Fluids B

• Powder Diffraction

• Review of Scientific Instruments

∗The Astronomical Journal and Applied Physics Letters are not part of this compuscript program. See Sec. V for more information on the Astronomical Journal. The Applied Physics Letters may be added to the compuscript program in the future; its weekly nature will make it more challenging.

III. WHERE TO TURN FOR HELP

In this section are listed both AIP contacts and editorial office contacts. Authors can always contact AIP with any questions, technical or procedural, surrounding the compuscript program. However, answers may be obtained more quickly by contacting the editorial office in question, depending on what stage the manuscript is in and whether the question is procedural or technical (involving the use of REVTEX).

3-1 \begin{matrix}c \\ ◯ \end{matrix} 1992 American Institute of Physics

A. Prior to manuscript acceptance

1. Procedural questions

Each editorial office has internal procedures to meet its needs. At the time this guide went to press, procedures for the editorial offices were being documented in the interest of publishing specific instructions for the authors at a future time.

For now, procedures may change frequently; therefore, every effort should be made to address procedural questions that arise prior to acceptance (such as, will you be contacted when the REVTEX file is needed, can the file be sent via electronic mail, etc.) with the editorial office in question rather than with AIP production staff.

However, if author questions cannot be properly resolved by the editorial office, contact AIP as directed below.

Electronic Submissions. Some editorial offices may be equipped to receive files for review purposes. Please contact the editorial office in question to determine the office’s capability and procedures.

2. REVTEX Questions

Questions about REVTEX use are best directed to AIP staff. See Sec. III C below.

B. After Manuscript Acceptance

Contact AIP staff with any questions that arise after the manuscript has been accepted for publication.

C. AIP Contacts

There are two contacts at AIP for compuscript processing questions: Peggy Sutherland and Liz Belmont (both are at 500 Sunnyside Boulevard, Woodbury, NY 11797). Please put questions in writing whenever possible or use electronic mail.

Liz Belmont 516/576-2454 e-mail: liz@aip.org

Peggy Sutherland 516/576-2335 e-mail: peggys@aip.org

D. Editorial Office Contacts

Editorial office contacts are listed below. In the pages that follow, the journal editor is listed first, and followed in parentheses by the office staff member (if available) who is able to answer procedural questions. Astronomical Journal^{∗}

Dr. Paul Hodge (Chaim Rosemarin) Dept. of Astronomy, FM-20 Univ. of Washington 206/685-2150 ASTRO@phast.phys.washington.edu ∗Do not use REVTEX. See Sec. V.

American Journal of Physics

Dr. Robert Romer (Karla Keyes) Merrill Science Bldg., Rm 222 P.O. Box 2262 Amherst College Amherst, MA 01002 413/542-5792

Chaos

Dr. David Campbell (Janis Bennett, AIP; 516/576-2403) Univ. of Illinois, Urbana-Champaign Loomis Lab. of Physics 1110 W. Green Street Urbana, IL 61801 217/333-3760 dkc@faust.physics.uiuc.edu

Computers in Physics

Dr. Lewis Holmes (Param Bajwa: 202/745-1895) AIP 1630 Connecticut Ave. NW Washington, DC 20009

Journal of Applied Physics

Dr. Steven Rothman (Cathey Dial) Argonne Natl. Lab. P.O. Box 8296 Argonne, IL 60439-8296 708/252-4200

The Journal of the Acoustical Society of America

Dr. Daniel Martin [Peggy Sutherland, AIP; 516/576-2335] 7349 Clough Pike Cincinnati, OH 45244

3-2 The Journal of Chemical Physics

Dr. John C. Light (Mitty Collier) James Franck Inst. Univ. of Chicago 5735 S. Ellis Avenue Chicago, IL 60637 312/702-7067

Journal of Mathematical Physics

Dr. Roger G. Newton (Hank Davis, 812/855-3576) Indiana University The Poplars, Rm 324 Bloomington, IN 47405 812/855-2095

Journal of Rheology

Dr. Arthur Metzner (Debbie Cook, x2299) Center for Composite Materials Univ. of Delaware Newark, DE 19716 302/831-2328

Journal of Vacuum Science and Technology A

Dr. Gerald Lucovsky (Becky York) Dept. of Physics North Carolina State Univ. Raleigh, NC 27650 919/248-1861

Journal of Vacuum Science and Technology B

Dr. Gary McGuire (Becky York) Microelectronics Center of North Carolina P.O. Box 12889 Research Triangle Park, NC 27709 919/248-1861

Medical Physics

Dr. John S. Laughlin (Linda Addonisio, 212/639-7414) Memorial Sloan-Kettering Cancer Center Medical Physics P.O. Box 62 1275 York Ave. New York, NY 10021

Publications of the Astronomical Society of the Pacific

Dr. Howard Bond (Denise Dankert) Space Telescope Science Institute 3700 San Martin Dr. Baltimore, MD 21218 410/338-4958 Physics of Fluids A

Dr. Andreas Acrivos (Tammy Erickson)

The Levich Inst., Steinman 202

City College of New York Convent Ave. at 140th St. New York, NY 10031 212/283-0962

Physics of Fluids B

Dr. Ronald C. Davidson (Ellen Webster)

Plasma Physics Lab. James Forrestal Campus Princeton Univ. P.O. Box 451 Princeton, NJ 08543 609/243-2425

Powder Diffraction

Deane K. Smith Dept. of Geosciences Pennsylvania State Univ. 239 Deike Bldg. University Park, PA 16802 815/865-5782

Review of Scientific Instruments

Dr. Thomas H. Braid

Argonne National Laboratory

P.O. Box 8293

Argonne, IL 60439-8293

708/252-8236

IV. OTHER PACKAGES THAT CAN BE USED WITH REVTEX

BIBTEX can be used to produce the bibliography section. Use the guidelines for Physical Review.

PostScript. If PostScript figures are available, provide them in files separate from the REVTEX file. Label the files FIG1, FIG2, FIG3a, etc. PostScript files may or may not be used for the production process. Glossy, cameraready figures must be provided to the editorial office for production even if PostScript figures are available.

V. BASIC INFORMATION

For most journals, the guidelines issued for APS and OSA compuscripts are acceptable. The exceptions are as follows:

• For the Astronomical Journal: Do not use REVTEX. The AAS is developing a macro package for use at this journal’s editorial office. AIP

3-3 will announce a compuscript program for AJ articles based on the AAS macros when software development has been completed.

• If the journal requires the author-year format for reference citations in text, as in the Journal of Rheology, do not use cite in the text. Type the reference explicitly within the text, and type the references in alphabetical order within the reference section. For example, in text: ... is the relaxation spectrum of the Rouse chain model [Bloomfield and Zimm (1966)]...

• Macro advice: Type-1 macros (those created by the author exclusively to save keystrokes) are discouraged. Although some macros may be acceptable in “reasonable” numbers and when simply constructed, authors may find that their compuscript has been rejected from the author-prepared program if macros cause complications in processing.

• More macro advice: Any time author-created macros are included in the file, the compuscript will require special handling, therefore reducing the time saved in production. Use of macros by authors will necessarily affect the feasibility of reducing page charges and of accelerated production schedules for compuscripts. Also, we would recommend strongly that author proofs are carefully proofread to ensure that macros have been properly processed.

• More macro advice: If we cannot process a file due to problem macros, the manuscript will be prepared conventionally to avoid production delays.

• Even more macro advice: We will not process author macros that give instructions to the TEX program (those that change internal counters, for example) or that involve recursive math definitions. If in doubt, remove macros before providing the file.

• Aligning math: Do not attempt to break and align math. The Xyvision composition system will automatically break and align displayed equations as needed.

VI. HOW TO USE REVTEX

AIP’s translation software will handle any compuscript that is formatted by following the REVTEX guidelines provided by the American Physical Society or the Optical Society of America. Review the manuscript samples provided by APS and OSA. Select a manuscript format based on your requirements; be sure to review Basic Information (Sec. V) above to ensure that your format choice is applicable.

In future releases of this guide, we may be able to recommend guidelines based on the AIP journal where the submission is being made. At this time, however, be assured that any REVTEX style selected will be supported by the AIP.

3-4